An accelerometer has the following properties:

Stiffness k = 500,000 N/m

Mass m = .01 Kg

Damping coefficient c = 95 N-sec/m

What is the damping ratio of the accelerometer?


1
Expert's answer
2021-03-18T03:36:13-0400

The damping ratio of the accelerometer can be found as


"\zeta=\frac{c}{2\sqrt{km}},"

where c - damping stiffness coefficient, k - stiffness, m - mass:

"\zeta=\frac{95}{2\sqrt{5·10^5·0.01}}=0.67."